OpenStudy (anonymous):

Sharon needs 3 3/8 yards of fabric to make a dress for herself and 2 1/8 yards of the same fabric to make a matching for her daughter. If she buys 6 1/8 yards of fabric , how much will she have left over?

OpenStudy (anonymous):

\[6\frac{ 1 }{ 8 } -( 3\frac{ 3 }{ 8 } + 2\frac{ 1 }{ 8 })\]

OpenStudy (anonymous):

can you solve that out?